Dan left downtown greensboro, and three hours later, Sara left going 51 mph faster to catch up. after another two hours, Sara caught up. Find dan's average speed
Answer by ankor@dixie-net.com(22740)

Dan left downtown greensboro, and three hours later, Sara left going 51 mph faster to catch up. after another two hours, Sara caught up.
Find dan's average speed
:
let s = Dan's speed
then
(s+51) = Sara's speed
:
From the information given, we know that S's travel time is 2 hrs, D's is 5 hrs
:
When S catches D, they will have traveled the same distance
Write a distance equation; dist = time * speed
5s = 2(s+51)
5s = 2s + 102
5s - 2s = 102
3s = 102
s = 102/3
s = 34 mph is D's speed
:
:
Confirm this, find the distance each traveled, should be equal
5 * 34 = 170 mi
2 * 85 = 170 mi